# The cohomological Kudla conjecture for unitary Shimura varieties

### v1

@@ 3.3. Toroidal compactification
The statements from this section follow closely the reference [How15, Section 3.3]. We assume that $\Gamma$ is a neat arithmetic subgroup of ${\mathbf{G}}_{1}(\mathbb{Q})$, and we refer to 3.7 for a discussion on how to reduce to this case.
@@
$X_{\Gamma}^{BB}=\Gamma_{L}\backslash\left(\mathbb{D}(V)\cup\bigcup_{\mathfrak{J}\subset L}\{P_{\mathfrak{J}}\}\right)\leavevmode\nobreak\ .$
@@
(3.1) $L=\mathfrak{J}\oplus\Lambda\oplus\mathfrak{c}\leavevmode\nobreak\ ,$
@@
$1\to N({\mathbf{P}})\to{\mathbf{P}}\to k^{\times}\times\mathrm{U}(\mathfrak{J}_{k}^{\perp}/\mathfrak{J}_{k})\to 1$
@@
$1\to\Gamma\cap N({\mathbf{P}})\to\Gamma\cap{\mathbf{P}}\to 1$
@@
$1\rightarrow C\rightarrow N({\mathbf{P}})\rightarrow W\rightarrow 1\leavevmode\nobreak\ .$
@@
for a diagonal matrix $A\in\mathrm{M}_{n}(\mathbb{Q})$ with positive diagonal entries and $\delta_{k}=i\sqrt{d_{k}}$, where $-d_{k}$ is the discriminant of $k$. The points of $N({\mathbf{P}})$ are given by:
$N({\mathbf{P}})(\mathbb{Q})=\left\{\begin{pmatrix}1&{}^{t}T&X\\ &\mathrm{I}_{n}&S\\ &&1\\ \end{pmatrix}\middle|\,\begin{array}[]{cc}T,S\in k^{n},&\delta_{k}T=-A\overline{S},\\ X\in k,&\delta_{k}(X-\overline{X})+^{t}SA\overline{S}=0\end{array}\right\}\leavevmode\nobreak\ .$
@@
$\Gamma\cap C(\mathbb{Q})=\left\{\begin{pmatrix}1&0&X\\ &\mathrm{I}_{n}&0\\ &&1\\ \end{pmatrix}|\,X\in r_{\mathfrak{J}}\mathbb{Z}\right\}\leavevmode\nobreak\ ,$
@@
The normal bundle of each boundary component ${\mathcal{B}}_{\mathfrak{J}}\subset X_{\Gamma}^{\mathrm{tor}}$ in the toroidal compactification pulls back to an anti-ample line bundle on $E_{M}$ whose class is a negative multiple of the polarization give by the positive definite quadratic form on $M$, viewed as a free $\mathbb{Z}$-module.
@@
If we view $\mathbb{D}(V)$ as an open subset of the complex projective space $\mathbb{P}(V_{\mathbb{C}})$, then linear projection from the point $[\mathfrak{J}_{\mathbb{C}}]$ gives a surjective morphism
@@
(3.2) $0\to\mathbb{Z}\to\Gamma\cap N({\mathbf{P}})\to M\to 0$
@@
One can check that $T_{\lambda}\in N({\mathbf{P}})$, and that it only depends on $\lambda$ via its equivalence class $[\lambda]\in\mathfrak{J}_{k}^{\perp}/\mathbb{Q}e$. Furthermore, for purely imaginary multiples of $e$ we have:
@@
Choosing any right inverse for the quotient map $\mathfrak{J}_{k}^{\perp}/\mathbb{Q}e\to\mathfrak{J}_{k}^{\perp}/\mathfrak{J}_{k}$, the formula $T_{\lambda}$ then gives a set-theoretic section of 3.2. The failure of $[\lambda]\mapsto T_{\lambda}$ to be a group homomorphism is measured by the following 2-cocycle for $M$:
@@ 3.4. Splitting of homology
$-\cap c_{1}(\mathcal{N}_{{\mathcal{B}}_{\mathfrak{J}}}):\mathrm{H}_{2g}(\mathcal{N}_{{\mathcal{B}}_{\mathfrak{J}}},\mathbb{Q})\to\mathrm{H}_{2g-2}(\mathcal{N}_{{\mathcal{B}}_{\mathfrak{J}}},\mathbb{Q}).$
Since the line bundle $N^{\vee}_{{{\mathcal{B}}_{\mathfrak{J}}}}$ is ample by Proposition 3.2, this surjectivity follows from Hard Lefschetz when $2g\leq n$.
@@ 3.5. Special cycles in the boundary
The first homology group of $E_{M}$ is identified with $\mathrm{H}_{1}(E,\mathbb{Z})\otimes_{\mathcal{O}_{k}}M\simeq M$ and the latter admits a $\mathbb{Z}$-valued symplectic pairing which is $\frac{1}{\sqrt{d_{k}}}\Im h$ where $h$ is the Hermitian pairing on $M$. This pairing is a morphism of Hodge structures and hence defines a Hodge class $D\in\mathrm{H}^{2}(E_{M},\mathbb{Z})$, which is algebraic by Lefschetz $(1,1)$-theorem. By Proposition 3.2, it is also equal to the Chern class of the conormal bundle of ${\mathcal{B}}_{\mathfrak{J}}$ in $X_{\Gamma}^{\mathrm{tor}}$.
@@
$Z(\underline{\lambda})=Z(\lambda_{1},\ldots,\lambda_{g})=Z(\lambda_{1})\cap\cdots\cap Z(\lambda_{g})\leavevmode\nobreak\ .$
@@
    If $\lambda_{1},\ldots,\lambda_{g}$ are linearly dependent over ${\mathcal{O}}_{K}$, then the class $[Z(\underline{\lambda})]$ vanishes in $\mathrm{CH}^{g}(E_{M})$.
@@
We first prove $(1)$. If one of the $\lambda_{i}$ is zero, then clearly we can remove it and the intersection has codimension at most $g-1$. Therefore, we can assume that all the $\lambda_{i}$ are non-zero.
For $\lambda\in M^{\vee}$ non-zero, notice that $[Z(\lambda)]=[u_{\lambda}^{-1}(\{x\})]+[u_{\lambda}^{-1}(\{-x\})]$ in $\mathrm{Pic}(E_{M})$ for any $x\in E(\mathbb{C})$. If $x$ is non-zero, then $Z(\lambda)$ has empty intersection with both $u_{\lambda}^{-1}(\{x\})$ and $u_{\lambda}^{-1}(\{-x\})$, therefore
$[Z(\lambda)].[Z(\lambda)]=0\quad\textrm{in}\quad\mathrm{CH}^{2}(E_{M})$
Let $u_{1},u_{2}\in\mathcal{O}_{k}\backslash\{0\}$. Then
$Z(u_{1}\cdot\lambda)=\bigcup_{\underset{u_{1}\cdot y=0}{y\in E(\mathbb{C})}}u_{\lambda}^{-1}\left(\{y\}\right),\,\textrm{and}\quad Z(u_{2}\cdot\lambda)=\bigcup_{\underset{u_{2}\cdot y=0}{y\in E(\mathbb{C})}}u_{\lambda}^{-1}\left(\{y\}\right)\leavevmode\nobreak\ .$
@@
$[Z(u_{2}\cdot\lambda)]=\sum_{\underset{u_{2}\cdot y=0}{y\in E(\mathbb{C})}}\left[u_{\lambda}^{-1}\left(\{\frac{y}{2}+x\}\right)\right]+\left[u_{\lambda}^{-1}\left(\{\frac{y}{2}-x\}\right)\right]\leavevmode\nobreak\ ,$
@@
$u_{\lambda}^{-1}\left(\{y\}\right)\cap u_{\lambda}^{-1}(\{\frac{y^{\prime}}{2}+x\}=\emptyset\leavevmode\nobreak\ ,\textrm{and}\quad u_{\lambda}^{-1}(\{y\}\cap u_{\lambda}^{-1}(\{\frac{y^{\prime}}{2}-x\}=\emptyset\leavevmode\nobreak\ .$
Hence $Z(u_{1}\cdot\lambda)$ and $Z(u_{2}\cdot\lambda)$ are linearly equivalent to divisors with empty intersection, hence $[Z(u_{1}\cdot\lambda)].[Z(u_{2}\cdot\lambda)]=0$. This proves the result for $g=2$. In general, let $\lambda_{1},\ldots,\lambda_{g}$ be linearly dependent over ${\mathcal{O}}_{K}$. Then we can assume that $u_{1}\cdot\lambda_{g}=\sum_{i=1}^{g-1}u_{i}\cdot\lambda_{i}$ where $u_{1}\neq 0$ and let $v=\sum_{i=1}^{g-1}u_{i}\cdot\lambda_{i}\neq 0$. Clearly $Z(\lambda_{1},\ldots,\lambda_{g-1})\subset Z(v)$ and $[Z(v)].[Z(\lambda_{g})]=0$ by the previous discussion, hence $Z(\lambda_{1},\ldots,\lambda_{g})=[Z(\lambda_{g})].[Z(\lambda_{1},\ldots,\lambda_{g-1})]=0$ in $\mathrm{CH}^{g}(E_{M})$. This concludes the proof of $(1)$.
@@
(3.2) $Z_{\mathfrak{J}}(\underline{\nu},N)=\bigcup_{\underset{h(\underline{\lambda},\underline{\lambda})=N}{\underline{\lambda}\in\underline{\nu}+M^{g}}}Z(\lambda_{1},\ldots,\lambda_{g})\leavevmode\nobreak\ .$
@@
The intersection of the Zariski closure of the special cycle ${\mathcal{Z}}(\underline{\nu},N)$ with the boundary ${\mathcal{B}}_{\mathfrak{J}}$ is empty if $\underline{\nu}\notin H_{\mathfrak{J}}^{\bot}$, and otherwise it is equal to $Z_{\mathfrak{J}}(\overline{\nu},N)$, where $\underline{\nu}\in M^{\vee}/M$ is the image of $\underline{\nu}$ under the projection map $H_{\mathfrak{J}}^{\bot}\rightarrow M^{\vee}/M$.
@@
We conclude that the validity of Theorem 1.4 and Corollary 1.6 for $X^{\mathrm{tor}}_{\Gamma^{\prime}}$ implies its validity for $X^{\mathrm{tor}}_{\Gamma}$.
@@ 4. Non-holomorphic geometric theta lift
In this section, we will exhibit a family of non-holomorphic modular forms defined by algebraic cycles that appear in the boundary of unitary Shimura varieties of signature $(n+1,1)$. The case of orthogonal Shimura varieties will be treated in a future work [EGT25].
We keep the notations from previous sections: $k\hookrightarrow\mathbb{C}$ is an imaginary quadratic field with ring of integers $\mathcal{O}_{k}$ and let $E=\mathbb{C}/\mathcal{O}_{k}$. Let $(M,h)$ be a positive definite Hermitian lattice of rank $n$ over $\mathcal{O}_{k}$. The abelian variety $E_{M}:=E\otimes_{{\mathcal{O}}_{K}}M$ admits a family of special cycles that were introduced in Section 3.5.
Our goal in this section is to investigate the modularity properties of the following generating series:
$\Phi_{M}^{g}(q)=\sum_{\underline{\lambda}\in{M^{\vee}}^{g}}[Z(\underline{\lambda})]q^{h(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\in\mathrm{CH}^{g}(E_{M})\otimes\mathbb{C}[(M^{\vee}/M)^{g}]\llbracket q\rrbracket\leavevmode\nobreak\ .$
@@
$\Phi^{g}(q)=\sum_{\underset{N\in\mathrm{Herm}_{g}(k)_{>0}}{\underline{\nu}\in(M^{\vee}/M)^{g}}}[Z(\underline{\nu},N)]q^{N}\mathfrak{e}_{\underline{\nu}}\in\mathrm{CH}^{g}(E_{M})\otimes\mathbb{C}[(M^{\vee}/M)^{g}]\llbracket q\rrbracket\leavevmode\nobreak\ .$
By Lemma 3.5, it is the intersection of the generating series of Kudla-Millson with the boundary divisor ${\mathcal{B}}_{\mathfrak{J}}$. Notice that only positive definite matrices $N$ appear, since the restriction of $\mathbb{E}$ to the boundary is trivial.

### v2

@@ 3.3. Toroidal compactification
The statements from this section follow closely the reference [How15, Section 3.3]. We assume that $\Gamma$ is a neat arithmetic subgroup of ${\mathbf{G}}_{1}(\mathbb{Q})$, and we refer to ˜3.7 for a discussion on how to reduce to this case.
@@
$X_{\Gamma}^{BB}=\Gamma_{L}\backslash\left(\mathbb{D}(V)\cup\bigcup_{\mathfrak{J}\subset L}\{P_{\mathfrak{J}}\}\right)\penalty 10000\ .$
@@
(3.1) $L=\mathfrak{J}\oplus\Lambda\oplus\mathfrak{c}\penalty 10000\ ,$
@@
$1\to N({\mathbf{P}})\to{\mathbf{P}}\to k^{\times}\times\mathrm{U}(\mathfrak{J}_{k}^{\perp}/\mathfrak{J}_{k})\to 1.$
@@
$1\to\Gamma\cap N({\mathbf{P}})\to\Gamma\cap{\mathbf{P}}\to 1.$
@@
$1\rightarrow C\rightarrow N({\mathbf{P}})\rightarrow W\rightarrow 1\penalty 10000\ .$
@@
for a diagonal matrix $A\in\mathrm{M}_{n}(\mathbb{Q})$ with positive diagonal entries and $\delta_{k}=i\sqrt{d_{k}}$, where $d_{k}$ is the discriminant of $k$. The points of $N({\mathbf{P}})$ are given by:
$N({\mathbf{P}})(\mathbb{Q})=\left\{\begin{pmatrix}1&{}^{t}T&X\\ &\mathrm{I}_{n}&S\\ &&1\\ \end{pmatrix}\middle|\,\begin{array}[]{cc}T,S\in k^{n},&\delta_{k}T=-A\overline{S},\\ X\in k,&\delta_{k}(X-\overline{X})+^{t}SA\overline{S}=0\end{array}\right\}\penalty 10000\ .$
@@
$\Gamma\cap C(\mathbb{Q})=\left\{\begin{pmatrix}1&0&X\\ &\mathrm{I}_{n}&0\\ &&1\\ \end{pmatrix}|\,X\in r_{\mathfrak{J}}\mathbb{Z}\right\}\penalty 10000\ ,$
@@
The normal bundle of each boundary component ${\mathcal{B}}_{\mathfrak{J}}\subset X_{\Gamma}^{\mathrm{tor}}$ in the toroidal compactification pulls back to an anti-ample line bundle on $E_{M}$ whose class is a negative multiple of the polarization given by the positive definite quadratic form on $M$, viewed as a free $\mathbb{Z}$-module.
@@
If we view $\mathbb{D}(V)$ as an open subset of the complex projective space $\mathbb{P}(V_{\mathbb{C}})$, then linear projection from the point $\mathfrak{J}_{\mathbb{C}}]$ gives a surjective morphism
@@
(3.2) $0\to\mathbb{Z}\to\Gamma\cap N({\mathbf{P}})\to M\to 0.$
@@
One can check that $T_{\lambda}\in N({\mathbf{P}})$, and that it only depends on $\lambda$ via its equivalence class $\lambda]\in\mathfrak{J}_{k}^{\perp}/\mathbb{Q}e$. Furthermore, for purely imaginary multiples of $e$ we have:
@@
Choosing any right inverse for the quotient map $\mathfrak{J}_{k}^{\perp}/\mathbb{Q}e\to\mathfrak{J}_{k}^{\perp}/\mathfrak{J}_{k}$, the formula $T_{\lambda}$ then gives a set-theoretic section of 3.2. The failure of $\lambda]\mapsto T_{\lambda}$ to be a group homomorphism is measured by the following 2-cocycle for $M$:
@@ 3.4. Splitting of homology
$\cap c_{1}(\mathcal{N}_{{\mathcal{B}}_{\mathfrak{J}}}):\mathrm{H}_{2g}(\mathcal{N}_{{\mathcal{B}}_{\mathfrak{J}}},\mathbb{Q})\to\mathrm{H}_{2g-2}(\mathcal{N}_{{\mathcal{B}}_{\mathfrak{J}}},\mathbb{Q}).$
Since the line bundle $N^{\vee}_{{{\mathcal{B}}_{\mathfrak{J}}}}$ is ample by Proposition˜3.2, this surjectivity follows from Hard Lefschetz when $2g\leq n$.
@@ 3.5. Special cycles in the boundary
The first homology group of $E_{M}$ is identified with $\mathrm{H}_{1}(E,\mathbb{Z})\otimes_{\mathcal{O}_{k}}M\simeq M$ and the latter admits a $\mathbb{Z}$-valued symplectic pairing which is $\frac{1}{\sqrt{d_{k}}}\Im h$ where $h$ is the Hermitian pairing on $M$. This pairing is a morphism of Hodge structures and hence defines a Hodge class $D_{{\mathfrak{J}}}\in\mathrm{H}^{2}(E_{M},\mathbb{Z})$, which is algebraic by Lefschetz $(1,1)$-theorem. By Proposition˜3.2, it is also equal to the Chern class of the conormal bundle of ${\mathcal{B}}_{\mathfrak{J}}$ in $X_{\Gamma}^{\mathrm{tor}}$.
@@
$Z(\underline{\lambda})=Z(\lambda_{1},\ldots,\lambda_{g})=Z(\lambda_{1})\cap\cdots\cap Z(\lambda_{g})\penalty 10000\ .$
@@
    If $\lambda_{1},\ldots,\lambda_{g}$ are linearly dependent over ${\mathcal{O}}_{K}$, then the class $Z(\underline{\lambda})]$ vanishes in $\mathrm{CH}^{g}(E_{M})$.
@@
We first prove $(1)$. If one of the $\lambda_{i}$ is zero, then clearly we can remove it and the intersection has codimension at most g1g-1. Therefore, we can assume that all the $\lambda_{i}$ are non-zero.
For $\lambda\in M^{\vee}$ non-zero, notice that $Z(\lambda)]=[u_{\lambda}^{-1}(\{x\})]+[u_{\lambda}^{-1}(\{-x\})]$ in $\mathrm{Pic}(E_{M})$ for any $x\in E(\mathbb{C})$. If $x$ is non-zero, then $Z(\lambda)$ has empty intersection with both $u_{\lambda}^{-1}(\{x\})$ and $u_{\lambda}^{-1}(\{-x\})$, therefore
$Z(\lambda)].[Z(\lambda)]=0\quad\textrm{in}\quad\mathrm{CH}^{2}(E_{M})$
[…]
$Z(u_{1}\cdot\lambda)=\bigcup_{\underset{u_{1}\cdot y=0}{y\in E(\mathbb{C})}}u_{\lambda}^{-1}\left(\{y\}\right),\,\textrm{and}\quad Z(u_{2}\cdot\lambda)=\bigcup_{\underset{u_{2}\cdot y=0}{y\in E(\mathbb{C})}}u_{\lambda}^{-1}\left(\{y\}\right)\penalty 10000\ .$
@@
$Z(u_{2}\cdot\lambda)]=\sum_{\underset{u_{2}\cdot y=0}{y\in E(\mathbb{C})}}\left[u_{\lambda}^{-1}\left(\{\frac{y}{2}+x\}\right)\right]+\left[u_{\lambda}^{-1}\left(\{\frac{y}{2}-x\}\right)\right]\penalty 10000\ ,$
@@
$u_{\lambda}^{-1}\left(\{y\}\right)\cap u_{\lambda}^{-1}(\{\frac{y^{\prime}}{2}+x\}=\emptyset\penalty 10000\ ,\textrm{and}\quad u_{\lambda}^{-1}(\{y\}\cap u_{\lambda}^{-1}(\{\frac{y^{\prime}}{2}-x\}=\emptyset\penalty 10000\ .$
Hence $Z(u_{1}\cdot\lambda)$ and $Z(u_{2}\cdot\lambda)$ are linearly equivalent to divisors with empty intersection, hence $Z(u_{1}\cdot\lambda)].[Z(u_{2}\cdot\lambda)]=0$. This proves the result for $g=2$. In general, let $\lambda_{1},\ldots,\lambda_{g}$ be linearly dependent over ${\mathcal{O}}_{K}$. Then we can assume that $u_{1}\cdot\lambda_{g}=\sum_{i=1}^{g-1}u_{i}\cdot\lambda_{i}$ where $u_{1}\neq 0$ and let $v=\sum_{i=1}^{g-1}u_{i}\cdot\lambda_{i}\neq 0$. Clearly $Z(\lambda_{1},\ldots,\lambda_{g-1})\subset Z(v)$ and $Z(v)].[Z(\lambda_{g})]=0$ by the previous discussion, hence $Z(\lambda_{1},\ldots,\lambda_{g})=[Z(\lambda_{g})].[Z(\lambda_{1},\ldots,\lambda_{g-1})]=0$ in $\mathrm{CH}^{g}(E_{M})$. This concludes the proof of $(1)$.
@@
(3.2) $Z_{\mathfrak{J}}(\underline{\nu},N)=\bigcup_{\underset{h(\underline{\lambda},\underline{\lambda})=N}{\underline{\lambda}\in\underline{\nu}+M^{g}}}Z(\lambda_{1},\ldots,\lambda_{g})\penalty 10000\ .$
@@
The intersection of the Zariski closure of the special cycle ${\mathcal{Z}}(\underline{\nu},N)$ with the boundary ${\mathcal{B}}_{\mathfrak{J}}$ is empty if $\underline{\nu}\notin H_{\mathfrak{J}}^{\bot}$ or if $N$ is not positive definite, and otherwise it is equal to $Z_{\mathfrak{J}}(\overline{\nu},N)$, where $\underline{\nu}\in M^{\vee}/M$ is the image of $\underline{\nu}$ under the projection map $H_{\mathfrak{J}}^{\bot}\rightarrow M^{\vee}/M$.
@@
We conclude that the validity of Theorem˜1.4 and Corollary˜1.6 for $X^{\mathrm{tor}}_{\Gamma^{\prime}}$ implies its validity for $X^{\mathrm{tor}}_{\Gamma}$.
@@ 4. Non-holomorphic geometric theta lift
In this section, we will exhibit a family of non-holomorphic modular forms defined by algebraic cycles that appear in the boundary of unitary Shimura varieties of signature $(n+1,1)$. The case of orthogonal Shimura varieties will be treated in a future work [EGT26].
We keep the notations from previous sections: $k\hookrightarrow\mathbb{C}$ is an imaginary quadratic field with ring of integers $\mathcal{O}_{k}$ and let $E=\mathbb{C}/\mathcal{O}_{k}$. Let $(M,h)$ be a positive definite Hermitian lattice of rank $n$ over $\mathcal{O}_{k}$. The abelian variety $E_{M}:=E\otimes_{{\mathcal{O}}_{K}}M$ admits a family of special cycles that were introduced in Section˜3.5.
Our goal in this section is to investigate the modularity properties of the following generating series valued [^5] We expect this generating series to also be modular with values in $\mathrm{CH}^{g}$, but proving that is not necessary for the present purpose. in the cohomology of $E_{M}$:
$\Phi_{M}^{g}(q)=\sum_{\underline{\lambda}\in{M^{\vee}}^{g}}[Z(\underline{\lambda})]q^{h(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\in\mathrm{H}^{2g}(E_{M})\otimes\mathbb{C}[(M^{\vee}/M)^{g}]\llbracket q\rrbracket\penalty 10000\ .$
@@
$\Phi^{g}_{M}(q)=\sum_{\underset{N\in\mathrm{Herm}_{g}(k)_{>0}}{\underline{\nu}\in(M^{\vee}/M)^{g}}}[Z(\underline{\nu},N)]q^{N}\mathfrak{e}_{\underline{\nu}}\in\mathrm{H}^{2g}(E_{M})\otimes\mathbb{C}[(M^{\vee}/M)^{g}]\llbracket q\rrbracket\penalty 10000\ .$
By Lemma˜3.5, it is the restriction of the generating series of Kudla-Millson with the boundary divisor ${\mathcal{B}}_{\mathfrak{J}}$. Notice that only positive definite matrices $N$ appear, since the restriction of $\mathbb{E}$ to the boundary is trivial.